### v1

Sequential Contractsthanks: This project has been partially funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 866132), by an Amazon Research Award, by the NSF-BSF (grant number 2020788), and by a grant from TAU Center for AI and Data Science (TAD). Tomer Ezra is supported by the Harvard University Center of Mathematical Sciences and Applications.
Tomer Ezra Harvard University, USA. Email: tomer@cmsa.fas.harvard.edu    Michal Feldman Tel Aviv University, Israel. Email: mfeldman@tauex.tau.ac.il    Maya Schlesinger Tel Aviv University, Israel. Email: mayas1@mail.tau.ac.il
We study the principal-agent setting, where a principal delegates the execution of a costly project to an agent. In the classical model, the agent chooses an action among a set of available actions. Every action is associated with some cost, and leads to a stochastic outcome for the project. The agent’s action is hidden from the principal, who only observes the outcome. The principal incentivizes the agent through a payment scheme (a contract) that maps outcomes to payments, with the objective of finding the optimal contract — the contract maximizing the principal’s expected utility.
In this work, we introduce a sequential variant of the model, capturing many real-life settings, where the agent engages in multiple attempts, incurring the sum of costs of the actions taken and being compensated for the best realized outcome. We study the contract design problem in this new setting. We first observe that the agent’s problem — finding the sequential set of actions that maximizes his utility for a given contract — is equivalent to the well-known Pandora’s Box problem. With this insight at hand, we provide algorithms and hardness results for the (principal’s) contract design problem, under both independent and correlated actions. For independent actions, we show that the optimal linear contract can be computed in polynomial time. Furthermore, this result extends to the optimal arbitrary contract when the number of outcomes is a constant. For correlated actions we find that approximating the optimal contract within any constant ratio is NP-hard.
1 Introduction
Contract theory is a pillar in microeconomics, addressing the challenge of motivating agents to exert effort when their actions are not directly observable. This issue is explored through the principal-agent model of Holmström (1979) and Grossman and Hart (1983).
In this model, a principal delegates the execution of a project to an agent, who chooses an action from a set of $n$ actions, each associated with a given cost and a probability distribution over $m$ possible outcomes. Each one of the outcomes generates some reward to the principal. The principal cannot observe the action chosen by the agent, but only the obtained outcome. To motivate the agent to take the desired action, the principal designs a contract, namely a payment scheme $t:[m]\rightarrow\mathbb{R}_{\geq 0}$ that specifies some payment to the agent for every observed outcome. Given a contract, the agent chooses an action that maximizes his expected utility, calculated as the expected payment minus the cost.
The optimal contract problem is the problem of finding the contract that maximizes the principal’s expected utility, calculated as the expected reward from the incentivized action minus the expected payment. This problem is poly-time solvable using linear programming (Grossman and Hart, 1983).
Contract design has traditionally been studied within the realm of economics. However, recent years have seen a surge of interest from computer science. This movement has expanded the study of contract design to encompass various algorithmic and computational perspectives, delving into the complexities of real-life scenarios. Among key areas explored are combinatorial settings that involve multiple agents (Babaioff et al., 2006; Emek and Feldman, 2012; Dütting et al., 2023a) and the consideration of combinatorial actions (Dütting et al., 2021a, 2023b; Ezra et al., 2024). This expansion into computer science has enriched the field, introducing new methodologies and insights into contract design.
We consider a combinatorial model where, similarly to the model of Dütting et al. (2021a), the agent can choose multiple actions. However, unlike their model, where the agent takes a set of actions simultaneously, our model captures scenarios in which the agent can choose the actions in a sequential manner.
The sequential multi-action model.
We introduce a contract design model with sequential actions that captures dynamic situations where an agent performs multiple actions in a sequential order, potentially adjusting their approach based on the results of prior efforts. This model reflects a wide array of real-world situations. For instance, in the recruitment process, a company (the principal) aims to fill a vacancy and engages a recruiter (the agent) for this purpose. The recruiter may conduct interviews with several candidates before making a hiring decision. Similarly, a freelance data scientist might test different methodologies and models, presenting only the most effective solution to the client. A teacher may use various techniques to prepare students for an exam, a doctor could try different treatments for a patient, and a researcher may explore multiple questions before choosing one for publication. Many real-life scenarios involve agents making several attempts to achieve a goal, yet traditional contract models often fail to consider this dynamic aspect of agent behavior. Our paper seeks to address this gap by enhancing the understanding of contract design in such dynamic contexts.
In the sequential multi-action contract model, a principal delegates the execution of a project to an agent. There is a set of $n$ actions available to the agent, each associated with some cost. There are $m$ possible outcomes, each associated with a non-negative reward to the principal. Every action is also associated with a random variable denoting that action’s stochastic outcome. These variables may be either independent or correlated. We refer to these cases as independent and correlated actions, respectively.
In our sequential model, the principal initially offers a contract, specifying a payment for every potential outcome. Following this, the agent engages in a series of actions sequentially, pursuing multiple attempts and obtaining outcomes, until opting to cease further actions. Specifically, at every point in time, the agent chooses some action, incurs its cost, and observes the realized outcome. The agent then decide whether to continue with another attempt using a different action or to halt. Upon halting, the agent chooses one of the outcomes achieved during the process as the final project’s outcome, which in turn determines the principal’s reward and the agent’s payment. His utility is then calculated as the payment received from the principal minus the cumulative costs of the actions undertaken.
The optimal contract is the contract that maximizes the principal’s utility, given that the agent responds with a sequential strategy that maximizes his own utility. The principal’s utility is the difference between the expected reward of the project’s final outcome and the expected payment to the agent.
1.1 Our Results
We study the computational complexity of the optimal (or approximately-optimal) contract in settings with sequential actions. We consider the problem with respect to independent actions as well as correlated actions, and with respect to the class of linear contracts as well as arbitrary contracts. A linear contract is one where the principal pays the agent a fixed fraction of the reward, determined by a parameter $\alpha\in[0,1]$ that specifies this fraction. This class of contracts is widely utilized in practical contract settings.
The independent-action model.
Our first result for the independent-action model is a poly-time algorithm that finds the optimal linear contract.
Theorem (Theorem 3.1) In the independent action model, the optimal linear contract can be computed in polynomial (in $n$ and $m$) time.
For arbitrary contracts, we devise an algorithm that finds the optimal contract in polynomial time, when $m$ (the number of outcomes) is constant.
Theorem (Theorem 3.2) In the independent action model, the optimal arbitrary contract can be computed in polynomial time, in cases where $m$ is constant.
The complexity of the optimal arbitrary contract for an arbitrary number of outcomes remains an interesting open problem. Notably, in Appendix C we construct an instance that admits $n^{\Omega(m)}$ best responses by the agent. This gives some indication of hardness, but does not in itself rule out a polytime algorithm for the problem.
Finally, we quantify the loss in the principal’s utility due to the restriction to linear contracts. Namely, we consider the worst-case ratio between the principal’s utility in the optimal arbitrary contract and the optimal linear contract, and find that this ratio can grow linearly with $n$ (the number of actions), even in the case of $m=3$ outcome.
Theorem (Theorem 3.3) There exists an instance with $n$ independent actions and $m=3$ outcomes such that the worst-case ratio between the optimal principal’s utility in an arbitrary contract and in a linear contract is $\Omega(n)$.
We also establish an upper bound of $O(n^{2}m)$ on this ratio, leveraging our analysis of the optimal linear contract, combined with a result by Dütting et al. (2019).
The correlated-action model.
In the correlated-action model, we establish a hardness of approximation result, showing that even in the binary-outcome case (where a project can either fail or succeed) it is NP-hard to approximate the optimal contract within any constant. This hardness also rules out a poly-time approximation of the optimal linear contract, since in the binary-outcome case the optimal contract is linear.
Theorem (Theorem 4.1) In the correlated action model, it is NP-hard to approximate the optimal (linear or arbitrary) contract within any constant, even in the binary-outcome case.
1.2 Our Techniques
We first observe that the agent’s problem (finding a strategy that maximizes his utility given a contract designed by the principal) reduces to the Pandora’s Box problem (Weitzman, 1978).
Weitzman (1978) presents a simple description of any optimal strategy for the Pandora’s Box problem, in cases where the values are independently distributed. This is directly applicable to the agent’s strategy in the independent-action setting. Within the framework of the agent’s problem in our sequential contract model, Weitzman’s algorithm calculates, given a contract $t$, a reservation value for each action $i$, and the agent’s optimal strategy (one that maximizes his utility) entails executing actions in non-decreasing order of reservation values, halting when the highest payment across all outcomes revealed thus far exceeds (or possibly meets) the subsequent reservation value.
Our results for the independent-action model rely on analyzing the reservation value as a function of the contract suggested by the principal. In the case of a linear contract, one can think of the reservation value as a function of the linear contract’s parameter $\alpha$. We show that this function is a convex piecewise-linear function with at most $m$ segments, and deduce that the agent’s optimal strategy may change only polynomially many times. By identifying the transition values of $\alpha$ — values at which the agent’s optimal strategy changes — we get polynomially-many potential candidates for the optimal linear contract, from which we can choose the best one.
For arbitrary contracts, with a constant number of outcomes $m$, the situation is more involved. We think of a contract as a vector in $\mathbb{R}_{\geq 0}^{m}$ and construct a hyperplane arrangement of polynomial size, ensuring that within any of its faces the set of all strategies that maximize the agent’s utility remains unchanged. This leads us to conclude that the optimal contract is located at a vertex of the hyperplane arrangement. Utilizing the fact that the hyperplane arrangement is of polynomial size in a vector space of a constant dimension, we infer that it has polynomially-many vertices. By iterating over these vertices we can find the optimal contract.
In the correlated-action model, we leverage an observation by Chawla et al. (2020), showing that in the binary-outcome case of the correlated Pandora’s Box problem, the optimal strategy admits a simple form. This observation directly informs the agent’s optimal strategy in the binary-outcome, correlated-action, sequential contracts setting.
Using this insight, we show that in this binary-outcome contract setting (where the project may either succeed or fail), the complex multi-dimensional distribution of the actions’ stochastic outcomes can effectively be captured by a “correlated OR” set function $f$, which maps each set of actions to the probability that at least one of them leads to a project success. We then show an equivalence between coverage functions and “correlated OR” functions, and construct a reduction from an NP-hard promise problem applying to coverage functions, introduced in (Ezra et al., 2024).
1.3 Related Work
The combinatorial-action setting.
The closest model to our setting is the combinatorial-action model introduced by Dütting et al. (2021a). In this model the agent chooses some subset of an available set of actions, with his choice stochastically determining the project’s success or failure according to a function $f$ that maps any subset of actions to a probability of success. The main result in Dütting et al. (2021a) is a poly-time algorithm for the optimal (linear) contract when $f$ is gross-substitutes.
This model has been further studied by Vuong et al. (2023) and Dütting et al. (2023b), who present a poly-time algorithm for computing the optimal contract for any class of functions that admits an efficient algorithm for the agent’s best-response (i.e., compute an optimal subset of actions given a contract) and poly-many “breakpoints” in the agent’s best-response. A direct corollary of this result is a poly-time algorithm for computing the optimal contract when the success probability function is supermodular and the cost function is submodular. Ezra et al. (2024) provide hardness of approximation results for this setting. They show that when $f$ is submodular it is NP-hard to approximate the optimal contract within any constant, and when $f$ is XOS, it is NP-hard to obtain a better approximation than $n^{-1/2+\varepsilon}$ for any $\varepsilon>0$.
The Pandora’s Box problem.
A closely related problem to the agent’s problem is the Pandora’s Box problem, introduced by Weitzman (1978), describing optimal search with costly inspection. In the Pandora’s Box problem there are $n$ boxes each with a stochastic hidden value and an associated cost. The decision maker opens the boxes sequentially. Upon opening a box, the hidden value is revealed and the associated cost is incurred. The goal is to maximize the expected maximum value revealed minus the expected total cost. Weitzman (1978) showed a simple optimal strategy for this problem, when the hidden values are distributed independently.
Chawla et al. (2020) showed that in the Pandora’s Box problem with correlated values, the optimal non-adaptive strategy is NP-hard to approximate within any constant, even with fully adaptive strategies. For a modified version in which one seeks to minimize the expected value plus total cost they show a poly-time constant-factor approximation algorithm for the optimal partially-adaptive strategy. Further work on the correlated values model focuses on the minimization variant and advances our understanding of approximation algorithms for this setting (Chawla et al., 2023; Gergatsouli and Tzamos, 2023).
Other combinatorial contract settings.
Another combinatorial model of contracts is the multi-agent setting, where the principal interacts with multiple agents. This model was first introduced by Babaioff et al. (2006), who focused on the binary-action, binary-outcome scenario, where agents have their own individual outcomes and some boolean function maps these to the project’s outcome. Emek and Feldman (2012) study this model in the special case where the project succeeds when at least one agent succeeds. This model was later extended by Dütting et al. (2023a), who consider a model where the success of the project is determined by a function $f:{\{0,1\}}^{[n]}\rightarrow[0,1]$, mapping the agent’s actions to a probability of success. Castiglioni et al. (2023) study a multi-agent setting in which each agent has his own outcome, and the principal’s reward depends on all individual outcomes. Dasaratha et al. (2023) consider a multi-agent setting with graph-based reward functions and continuous effort levels.
Additional literature on combinatorial contracts includes settings with exponentially many outcomes (Dütting et al., 2021b) and settings with multiple principals (Alon et al., 2021).
2 Model and Preliminaries
We consider a principal-agent setting where the principal delegates the execution of a project to an agent. In order to execute the project, the agent can take actions in a sequential way from a set $A$ of $n$ possible actions, where each action $i\in A$ is associated with a cost $c_{i}\geq 0$. The execution of the project leads to an outcome in a set $[m]$ of $m$ possible outcomes, where each outcome $j\in[m]$ is associated with a reward for the principal $r(j)\geq 0$. We assume without loss of generality that $r(1)\leq\dots\leq r(m)$, and that $r(1)=0$. We sometimes refer to $1$ as the zero outcome.
Actions lead to outcomes stochastically. Specifically each action $i\in A$ is associated with a random variable $X_{i}$ that attains values within $[m]$, denoting that action’s outcome. The variables $\{X_{i}\}_{i\in A}$ are distributed according to some joint distribution $\mathcal{F}$, and are either independent or correlated. The case where they are independent is referred to as the independent action model and the case where they may be correlated is referred to as the correlated action model. If they are independent, then the distribution of each $X_{i}$ is given by a vector $\{p_{ij}\}_{j\in[m]}$, where $p_{ij}$ is the probability that $X_{i}=j$. If the actions are correlated, then for every realization vector $v\in[m]^{n}$, $p(v)$ denotes the probability that $v$ is realized.
The agent’s choices are hidden from the principal, who can observe only the project’s final outcome, for which she is rewarded. To incentivize the agent to take actions, the principal designs a contract $t:[m]\rightarrow\mathbb{R}_{\geq 0}$ which maps each outcome to a non-negative payment to the agent.
Sequential model.
In a sequential model, once a contract has been determined by the principal, the agent takes actions sequentially. At each point in time, the agent selects an action $i\in A$, incurs the associated cost $c_{i}$, and observes the outcome that is revealed as a result of this action (i.e., the value of $X_{i}$). The agent then decides whether to halt or continue with another action (that has not yet been taken). When the agent halts, he selects one outcome from those revealed, and this selected outcome becomes the final outcome of the project, upon which the agent is paid.
A strategy of the agent in this scenario, denoted $\pi$, determines which action to take at each point in time, and may choose to halt. When halting, the agent chooses which of the obtained outcomes is the final project’s outcome, for which the principal is rewarded and the agent is paid. If the agent decides to halt before taking any action, then the zero outcome is the final outcome, which we assume is always revealed. All decisions may depend on the instance, the contract, the actions taken so far, and their revealed outcomes.
Given an agent’s strategy $\pi$, we denote by $S(\pi)$ the random variable depicting the set of actions taken by the agent, and by $\omega(\pi)$ the final outcome.
The agent’s utility under a contract $t$ and a strategy $\pi$ is the expected payment from the principal minus the expected sum of the costs of the actions taken,
$u_{A}(t,\pi)=\mathbb{E}_{X_{1},\dots,X_{n}\sim\mathcal{F}}[t(\omega(\pi))-\sum_{i\in S(\pi)}c_{i}].$
The principal’s utility under a contract $t$ and a strategy $\pi$ is the expected reward minus the expected payment, i.e.,
$u_{P}(t,\pi)=\mathbb{E}_{X_{1},\dots,X_{n}\sim\mathcal{F}}[r(\omega(\pi))-t(\omega(\pi))].$
Best response.
The agent’s best response to a contract $t$, denoted $\pi^{\star}(t)$, is a strategy that maximizes the agent’s utility, namely $\pi^{\star}(t)\in\text{argmax}_{\pi}u_{A}(t,\pi)$. As standard in the literature, the agent breaks ties in favor of the principal’s utility. When both the agent and the principal are indifferent, ties are broken arbitrarily and consistently [^1] Since the set of strategies is finite, one can impose a full order over the set, thus consistently is well defined..
We let $u_{P}(t)$ denote the principal’s utility under the contract $t$ given that the agent best responds, i.e., $u_{P}(t)=u_{P}(t,\pi^{\star}(t))$. Given a family of contracts $\mathcal{T}$, the principal’s problem is to find a contract $t\in\mathcal{T}$ that maximizes $u_{P}(t)$.
Linear contracts.
A prevalent class of contracts is linear contracts, where the principal compensates the agent with a constant fraction of the reward. A linear contract is given by a parameter $\alpha\in[0,1]$, where the payment is $t(j)=\alpha\cdot r(j)$. We denote a linear contract with parameter $\alpha$ by $t_{\alpha}$.
Approximation algorithms.
A $\beta$-approximation algorithm for the optimal contract returns a contract $t$ such that $u_{P}(t)\geq\beta\cdot u_{P}(t^{\star})$, where $t^{\star}$ is an optimal contract.
2.1 Pandora’s Box Problem
A closely related problem is the well known Pandora’s Box problem, introduced in (Weitzman, 1978). This problem models a search for the best alternative under uncertainty, when inspections are costly. In the Pandora’s Box problem there is a set $B$ of $n$ boxes. Each box $i\in B$ is associated with an inspection cost $c_{i}\geq 0$ and a random value $V_{i}$. The variables $\{V_{i}\}_{i\in B}$ are distributed according to a joint distribution $\mathcal{F}$.
At each point in time, a decision maker selects a box $i\in B$ to open, incurring cost $c_{i}$, and revealing the value hidden inside (the value of the random variable $V_{i}$). The decision maker then decides whether to halt, in which case he obtains the maximum value revealed until this point, or continue and open another (yet-unopened) box.
A strategy for a problem instance of Pandora’s Box defines which box to open next and when to halt. For a given strategy $\pi$, let $S(\pi)$ denote the (random) set of boxes opened, and $V(\pi)=\max_{i\in S(\pi)}V_{i}$ denote the (random) maximum value observed by $\pi$, this is interpreted as 00 if the strategy opens no boxes. The strategy’s expected utility is given by
$u(\pi)=\mathbb{E}_{\{V_{i}\}_{i\in B}\sim\mathcal{F}}[V(\pi)-\sum_{i\in S(\pi)}c_{i}].$
Relationship between the agent’s problem and Pandora’s Box.
The agent’s problem in our sequential model, i.e., finding the best response $\pi^{\star}(t)$ to a contract $t$, relates very naturally to an instance of the Pandora’s Box problem. Indeed, given a contract $t$, to maximize his utility, the agent seeks to solve the Pandora’s Box problem instance in which the set of boxes corresponds to the set of actions, the cost of box $i$ equals the cost of action $i$, and the value of box $i$ is $V_{i}=t(X_{i})$, i.e., the payment for the outcome of the action associated with box $i$. We call this instance of the Pandora’s Box problem the induced Pandora’s Box instance. Note that solving the induced Pandora’s Box instance doesn’t necessarily solve the problem of the agent’s best response since, when more than one strategy maximizes the agent’s utility, it may not be trivial which of them leads to the best principal’s utility.
3 Independent Actions
In this section, we consider the independent action model.
We begin our analysis, in Section 3.1, by characterizing the agent’s optimal strategies (strategies which maximize his utility) under a given contract in the independent-action setting. This characterization gives insight into how these optimal strategies change when the contract changes. Additionally, we show that it is possible to compute the principal’s utility from a contract in polynomial time.
In Section 3.2, we present our first result: a polytime algorithm for the optimal linear contract. In Section 3.3 we consider the problem of computing the best arbitrary contract and show that this can be done in polynomial time in cases where $m$ is constant. Finally, in Section 3.4 we quantify the worst-case ratio between the principal’s utility in an arbitrary and linear contract, and show that this ratio can be as high as $\Omega(n)$ but is upper bounded by $O(n^{2}m)$.
3.1 The Agent’s Optimal Strategies
We call any strategy $\pi$ that maximizes the agent’s utility in response to a contract $t$ an optimal strategy. As noted in Section 2, the agent’s optimal strategies are the optimal strategies in the induced Pandora’s Box instance. Weitzman (1978) gives a simple description of these strategies in the independent-action model.
Weitzman’s algorithm.
Weitzman (1978) gives a simple description of all optimal strategies in a Pandora’s Box problem instance where the variables $\{V_{i}\}_{i\in B}$ are independent (i.e., for every $i\in B$, $V_{i}$ is sampled independently from a probability distribution $\mathcal{F}_{i}$). At the core of Weitzman’s algorithm is the notion of a reservation value of box $i$, defined as the value $z_{i}$ satisfying
$\mathbb{E}_{\{V_{i}\}_{i\in B}\sim\mathcal{F}}[(V_{i}-z_{i})^{+}]=c_{i},$ (1)
where $(X)^{+}$ denotes $\max(X,0)$. When $c_{i}=0$, we define $z_{i}=\infty$.
Weitzman (1978) shows that any optimal strategy opens the boxes in non-increasing order of $z_{i}$, until the maximum value revealed so far [^2] We interpret the maximum revealed value to be 00 if no boxes have been opened yet. exceeds (or, possibly, meets) the next reservation value, or until all boxes have been opened. Furthermore, any strategy satisfying these conditions is optimal. More formally,
Proposition 3.1 (Weitzman (1978)).
A strategy $\pi$ of the Pandora’s Box problem is optimal if and only if the following conditions are always satisfied at each point in time:
  1. 1.
    When $\pi$ opens box $i\in B$, it holds that $z_{i}\geq z_{i^{\prime}}$ for any unopened box $i^{\prime}\in B$.
  2. 2.
    Let $v$ be the maximum value revealed so far, and let $i\in B$ be an unopened box of maximum $z_{i}$. If $v>z_{i}$, $\pi$ halts, and if $v<z_{i}$, $\pi$ doesn’t halt.
This implies that an agent’s strategy is optimal if and only if it satisfies the conditions of Proposition 3.1 in the induced Pandora’s Box instance and upon halting chooses a revealed outcome that maximizes the agent’s payment. More formally, for any action $i\in A$, we denote by $z_{i}(t)$ the reservation value of box $i$ in the Pandora’s Box instance induced by a contract $t$, and get the following:
Observation 3.1.
For any contract $t$, an agent’s strategy $\pi$ is optimal if and only if the following conditions are always satisfied at each point in time:
  1. 1.
    When $\pi$ takes an action $i\in A$ it holds that $z_{i}(t)\geq z_{i^{\prime}}(t)$ for any untaken action $i^{\prime}\in A$.
  2. 2.
    Let $j\in[m]$ be a revealed outcome of maximum payment, and let $i\in A$ be an untaken action of maximum $z_{i}(t)$. If $t(j)>z_{i}(t)$, $\pi$ halts, and if $t(j)<z_{i}(t)$, $\pi$ doesn’t halt.
  3. 3.
    Upon halting, it holds that $t(\omega(\pi))\geq t(j)$ for any revealed outcome $j$ (recall that $\omega(\pi)$ is the final outcome chosen by $\pi$, see Section 2).
Computing the principal’s utility.
We now turn to the task of computing the principal’s utility from a given contract. Specifically, we show the following proposition.
Proposition 3.2.
The principal’s utility from a contract is computable in polynomial time.
There are a few challenges we face here. Firstly, the number of optimal strategies may be exponential, so finding an optimal strategy that also maximizes the principal’s utility is non-trivial. Also, despite the existence of very simple optimal strategies, some optimal strategies may be complex, and it is not clear if the principal’s utility from a complex strategy can be computed efficiently.
To address these challenges, for any contract $t$ and $\varepsilon>0$, we define the contract $t^{\varepsilon}$ as $t^{\varepsilon}(j)=t(j)+\varepsilon\cdot(r(j)-t(j))$. We then show that for a small enough $\varepsilon$, any optimal strategy under the contract $t^{\varepsilon}$, is also an optimal strategy under $t$. We continue to show that in such a case, an optimal strategy under the contract $t^{\varepsilon}$ also maximizes the principal’s utility under $t$.
We then restrict our attention to a simple type of agent’s strategies which we call non-adaptive strategies. In such strategies, the agent’s choice of the order in which to take actions is fixed in advance, and his choice of when to halt depends only on the maximum revealed outcome so far. It is easy to see that non-adaptive optimal strategies exist (and are easy to find). The important property of non-adaptive strategies is that the principal’s utility when the agent engages in one can be computed efficiently. Therefore, by finding a strategy $\pi$ that is a non-adaptive optimal strategy under the contract $t^{\varepsilon}$, we can compute $u_{P}(t)$ by computing $u_{P}(t,\pi)$.
For the full details, see Appendix A.
3.2 Optimal Linear Contracts
In this section, we prove that the optimal linear contract can be found in polynomial time. More formally, we prove the following theorem:
Theorem 3.1.
In the independent action model, the optimal linear contract can be computed in polynomial (in $n$ and $m$) time.
Our proof of Theorem 3.1 relies on the concept of critical values, so we begin by introducing it.
Critical values.
The concept of critical values follows from an analysis by Dütting et al. (2019). Put in terms of our problem, for any strategy $\pi$ we denote by $c(\pi)=\mathbb{E}[\sum_{i\in S(\pi)}c_{i}]$ the strategy’s expected cost to the agent, and by $R(\pi)=\mathbb{E}[r(\omega(\pi))]$ the strategy’s expected principal’s reward. Under a linear contract $t_{\alpha}$ and a strategy $\pi$, the agent and principal’s utilities are given by
$u_{A}(t_{\alpha},\pi)=\alpha\cdot R(\pi)-c(\pi)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ u_{P}(t_{\alpha},\pi)=(1-\alpha)R(\pi)$
From this description of the principal’s utility it follows that strategies that maximize $R(\pi)$ are the ones that maximize the principal’s utility. From the description of the agent’s utility it follows that the agent’s optimal strategies only changes when $\alpha=\frac{c(\pi_{1})-c(\pi_{2})}{R(\pi_{1})-R(\pi_{2})}$ for some two strategies $\pi_{1},\pi_{2}$ with $R(\pi_{1})\neq R(\pi_{2})$. At such points of $\alpha$, the agent is indifferent between the strategies $\pi_{1}$ and $\pi_{2}$ (and possibly other strategies), and therefore breaks ties in favor of the highest expected reward, which, as one can verify, are the same strategies that maximize his utility to the right of $\alpha$. Due to consistent tie-breaking we get that agent’s best response $\pi^{\star}(t_{\alpha})$ only changes finitely many times and is right-continuous.
The values of $\alpha$ where the agent’s best response changes are called the critical values. An important note to stress that follows from this analysis is the following:
Observation 3.2.
If $\alpha$ is a critical value then $\pi^{\star}(t_{\alpha})$ — the agent’s best response to the contract $t_{\alpha}$ — wasn’t an agent’s optimal strategy under any linear contract $t_{\alpha^{\prime}}$ where $\alpha^{\prime}<\alpha$.
It is easy to see that the optimal value of $\alpha$ is also a critical value since the principal seeks to minimize $\alpha$ while incentivizing the agent to engage in some strategy. By enumerating over the critical values and computing the principal’s utility for each, we can choose the best among them and are guaranteed it maximizes the principal’s utility.
Therefore, to prove Theorem 3.1, it suffices show that iterating over all critical values can be done in polynomial time (in $n$ and $m$). This is established in the following proposition.
Proposition 3.3.
In the independent action model, there are $O(n^{2}m)$ critical values of $\alpha$, and they are computable in polynomial time.
Proof.
To prove Proposition 3.3 we characterize $z_{i}(\alpha)$ — the reservation values as a function of $\alpha$ — and show a relationship between $z_{i}(\alpha)$ and the critical values of $\alpha$.
The characterization of $z_{i}(\alpha)$ is cast in the following lemma.
Lemma 3.1.
For every action $i\in A$, $z_{i}(\alpha)$ is a convex piecewise linear function with at most $m$ segments and is given by an efficiently computable closed-form expression.
The proof of Lemma 3.1 is deferred to the appendix (see Appendix B.1).
We now describe a relation between the values of $z_{i}(\alpha)$, and the critical points of $\alpha$.
Lemma 3.2.
If $\alpha^{\prime}$ is a critical value, then one of the following holds:
  1. 1.
    For some two actions $i_{1}\neq i_{2}$, it holds that $z_{i_{1}}(\alpha)$ and $z_{i_{2}}(\alpha)$ intersect at $\alpha^{\prime}$.
  2. 2.
    For some action $i\in A$ and outcome $j\in[m]$, it holds that $z_{i_{1}}(\alpha)$ intersects with $\alpha\cdot r(j)$ at $\alpha^{\prime}$.
Lemma 3.2 follows from an analysis of the conditions in Observation 3.1. The full details of the proof can be found in Appendix B.2.
Overall, we have that only points which that satisfy either condition of Lemma 3.2 may be critical values. Any two piecewise linear functions with at most $m$ segments intersect at most $m$ times, which, due to Lemma 3.1, implies that at most $O(n^{2}m)$ points satisfy the first condition of Lemma 3.2. Since $z_{i}(\alpha)$ is convex, it intersects with any straight line at most twice, which implies at most $O(n\cdot m)$ points satisfy the second condition of Lemma 3.2. The values of $\alpha$ at which these intersections happen can be computed from our explicit expression for $z_{i}(\alpha)$ (given by Lemma 3.1), and are the only candidates for critical values of $\alpha$, which concludes our proof. ∎
3.3 General Contracts
In this section, we prove that when $m$ is constant, the best monotone contract can also be found in polynomial time. This is cast in the following theorem.
Theorem 3.2.
In the independent action model, the optimal arbitrary contract can be computed in polynomial time, in cases where $m$ is constant.
Remark 3.1.
In Appendix C, we show that the number of best responses by the agent to general contracts may be as large as $n^{\Omega(m)}$. Furthermore, this lower bound also holds for the family of monotone contracts. This suggests an inherent complexity of the optimal contract problem, but doesn’t rule out the existence of a polytime algorithm for it.
Before turning to the proof of Theorem 3.2, we set up a framework. Firstly, we expand our analysis from Section 3.1 and examine more closely how the agent’s optimal strategies change as a function of the contract. Then, as our proof of Theorem 3.2 relies heavily on hyperplane arrangements, we provide some preliminaries regarding these.
The set of optimal strategies.
Based on the description of the agent’s optimal strategies given in Observation 3.1, all optimal strategies under some contract $t$ are determined by a 4-tuple $(P_{A},P_{[m]},\tau^{>},\tau^{<})$ where
  • $P_{A}$ is the full preorder [^3] A preorder $P$ is a reflexive and transitive binary relation. A full preorder also satisfies $x\leq_{P}y$ or $y\leq_{P}x$ for any $x,y$. induced on the set of actions $A$ according to $i_{1}\leq_{P_{A}}i_{2}$ if and only if $z_{i_{1}}(t)\leq z_{i_{2}}(t)$. $P_{A}$ determines the possible action-taking order of an optimal strategy.
  • $P_{[m]}$ is the full preorder induced on the set of outcomes $[m]$ according to $j_{1}\leq_{P_{[m]}}j_{2}$ if and only if $t(j_{1})\leq t(j_{2})$. $P_{[m]}$ determines which outcomes an optimal strategy may choose upon halting.
  • $\tau^{>}=\{\tau^{>}_{i}\}_{i\in A}\subseteq 2^{[m]}$ and $\tau^{<}=\{\tau^{<}_{i}\}_{i\in A}\subseteq 2^{[m]}$ are defined by
    $\tau^{>}_{i}=\{j\in[m]\mid t(j)>z_{i}(t)\}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \tau^{<}_{i}=\{j\in[m]\mid t(j)<z_{i}(t)\}.$
    $\tau^{>}$ and $\tau^{<}$ determine when an optimal strategy must halt and when it must continue.
We sometimes use $P_{A}(t),P_{[m]}(t),\tau^{>}(t),\tau^{<}(t)$ denote the values of $P_{A},P_{[m]},\tau^{>},\tau^{<}$ under a contract $t$, but usually we neglect the contract which defines them for brevity.
Hyperplane arrangements.
A hyperplane arrangement is a set $\mathcal{A}$ of hyperplanes in some vector space. We focus on hyperplanes over the reals. Consider some hyperplane arrangement $\mathcal{A}$ over $\mathbb{R}^{d}$. When the hyperplanes of $\mathcal{A}$ are removed, the space is divided into components, each one being an open $d$-dimensional polytope. These are called the regions of $\mathcal{A}$.
A face of an arrangement $\mathcal{A}$ is an intersection of the closure of a region with some intersection of the hyperplanes in $\mathcal{A}$, i.e,
$\bar{R}\cap\bigcap_{A\in\mathcal{I}}A,$
where $R$ is a region of $\mathcal{A}$ and $\mathcal{I}\subseteq\mathcal{A}$.
The dimension of a face $F$ is defined as the dimension of the affine span of $F$. A $k$-face of $\mathcal{A}$ is a face of dimension $k$.
Two useful definitions in this context are the relative interior of a set $S$, which we denote by $\text{relint}(S)$ and is defined as the interior of $S$ within the affine span of $S$, and the relative boundary of $S$, which is its boundary within the affine space. An important property is that, for any $k>0$, the relative boundary of any $k$-face is comprised of $k-1$-faces.
Observation 3.3.
[…]
Finally, we rely on the following proposition:
Proposition 3.4 (Zaslavsky (1975)).
The number of $k$-faces in a hyperplane arrangement of size $n$ in $\mathbb{R}^{d}$ is at most
(${n\choose d-k}\sum_{i=0}^{k}{n-d+k\choose i}.$
This implies that the number of faces of an arrangement $\mathcal{A}$ is bounded by $O(|\mathcal{A}|^{d})$.
Proof of Theorem 3.2.
Denote $L=\max_{j\in[m],i\in A:p_{ij}\neq 0}\frac{r(m)}{p_{ij}}$. It is easy to verify that any optimal contract $t$ satisfies $t(j)\in[0,L]$ for any outcome $j\in[m]$. For our proof we think of a contract $t$ as an element of $[0,L]^{m}$, where the $j$th coordinate of the vector is the payment for outcome $j$. The basic idea is to construct a hyperplane arrangement which captures changes in the 4-tuple $(P_{[m]},P_{A},\tau^{>},\tau^{<})$. The intuition here is that, in the same way that under linear contracts the points where the agent’s best response changes are the candidates for the optimal contracts, in a general arbitrary contract the points where the 4-tuple changes are of interest. More formally, if within the relative interior of each face of a hyperplane arrangement the 4-tuple $(P_{[m]},P_{A},\tau^{>},\tau^{<})$ is constant, we show that the optimal contract is a 00-face (a point which is the intersection of $m$ hyperplanes).
We begin by characterizing the reservation value as piecewise linear. This is essential in order to capture changes in $P_{A}$ with hyperplanes.
Lemma 3.3.
For any $i\in A$, there exist $2^{m}$ linear functions, denoted by $z^{S}_{i}(t)$ for any $S\in 2^{[m]}$, such that for any contract $t$ it holds that $z_{i}(t)=z^{S}_{i}(t)$, where $S=\tau^{>}_{i}(t)$.
Proof.
Let $i\in A$. For any subset $S\subseteq[m]$ we define $z^{S}_{i}(t)$ as follows:
$z^{S}_{i}(t)=\frac{\sum_{j\in S}p_{ij}\cdot t(j)-c_{i}}{\sum_{j\in S}p_{ij}},$
Let $t$ be some contract. The random variable $t(X_{i})$ attains values within $\{t(j)\}_{j\in[m]}$ and is equal to $t(j)$ with probability $p_{ij}$. Plugging this into Equation (1) we get
$c_{i}=\mathbb{E}[(t(X_{i})-z_{i}(t))^{+}]=\sum_{j\in[m]:t(j)>z_{i}(t)}p_{ij}(t(j)-z_{i}(t))=\sum_{j\in\tau^{>}_{i}(t)}p_{ij}(t(j)-z_{i}(t)),$
where the last equality is by the definition of $\tau^{>}_{i}(t)$. This simplifies into
$z_{i}(t)=\frac{\sum_{j\in\tau^{>}_{i}(t)}p_{ij}\cdot t(j)-c_{i}}{\sum_{j\in\tau^{>}_{i}(t)}p_{ij}}=z^{\tau^{>}_{i}(t)}_{i}(t),$
as needed. ∎
We now begin to formalize our earlier notion of the candidates for the optimal contract being 00-faces of an arrangement. We begin our analysis by considering the principal’s utility in polytopes (which, in a sense, are the basic building block of hyperplane arrangements).
Lemma 3.4.
Let $Q\subseteq[0,L]^{m}$ be a polytope. $u_{P}(t)$ attains a maximum [^4] The principal’s utility is not continuous therefore this is not immediate from the Extreme Value Theorem. on $Q$.
The proof of this lemma is deferred to Appendix B.3.
Furthermore, when the agent’s optimal strategies remain the same throughout the relative interior of a polytope, this maximum is attained at the boundary. More formally,
Lemma 3.5.
Let $Q\subseteq[0,L]^{m}$ be a polytope such that in the relative interior of $Q$ the 4-tuple $(P_{A},P_{[m]},\tau^{>},\tau^{<})$ is constant. The principal’s utility is maximized at the relative boundary of $Q$.
The proof of this lemma relies on two observations. Firstly, that for any single strategy $\pi$ the principal’s utility $u_{P}(t,\pi)$ is maximized at a vertex. Secondly, any strategy which maximizes the agent’s utility at the relative interior of a polytope also maximizes it at the boundary. The full details are deferred to Appendix B.4. As a corollary, we get the following key finding.
Corollary 3.1.
[…]
To prove Theorem 3.2, it suffices to show a hyperplane arrangement of polynomial size that can be computed in polynomial time and satisfies the conditions of the Corollary 3.1. This is because, in a vector space of constant dimension, all 00-faces of a hyperplane arrangement can be found in polytime.
Consider the following hyperplane arrangements:
  1. 1.
    $\mathcal{A}_{1}=\{\{t\in\mathbb{R}^{m}\mid t_{j}=0\}\mid j\in[m]\}\cup\{\{t\in\mathbb{R}^{m}\mid t_{j}=L\}\mid j\in[m]\}$
  2. 2.
    $\mathcal{A}_{2}=\{E_{j_{1},j_{2}}\mid j_{1}\neq j_{2}\in[m]\}$, where $E_{j_{1},j_{2}}$ is defined as
    $E_{j_{1},j_{2}}=\{t\in\mathbb{R}^{m}\mid t_{j_{1}}=t_{j_{2}}\}.$
  3. 3.
    $\mathcal{A}_{3}=\{T_{\rho,i,j}\mid\rho:[m]\rightarrow[m]\text{ is a bijection},i\in A,j\in[m]\}$, where $T_{\rho,i,j}$ is defined as
    $T_{\rho,i,j}=\{t\in\mathbb{R}^{m}\mid\sum_{k=j}^{m}p_{ik}(t(\rho(k))-t(\rho(j)))=c_{i}\}.$
  4. 4.
    $\mathcal{A}_{4}=\{Z_{i_{1},i_{2},S_{1},S_{2}}\mid i_{1}\neq i_{2}\in A,S_{1},S_{2}\subseteq[m]\}$ where $Z_{i_{1},i_{2},S_{1},S_{2}}$ is defined as
    $Z_{i_{1},i_{2},S_{1},S_{2}}=\{t\in\mathbb{R}^{m}\mid z^{S_{1}}_{i_{1}}(t)=z^{S_{2}}_{i_{2}}(t)\}.$
We consider the hyperplane arrangement $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}\cup\mathcal{A}_{3}\cup\mathcal{A}_{4}$, and show that it satisfies the conditions of Corollary 3.1.
It is immediate that $\mathcal{A}_{1}$ guarantees the condition that any face $F$ satisfies $F\subseteq[0,L]^{m}$ or $\text{relint}(F)\cap[0,L]^{m}=\emptyset$, i.e., it creates a separation between regions which define a contract ($[0,L]^{m}$) and regions which do not. The hyperplanes of $\mathcal{A}_{2}$ capture the transition points in $P_{[m]}$, these are points at which the agent’s preference between outcomes might change. More specifically, each hyperplane $E_{j_{1},j_{2}}$ captures the places where the agent may change his preference between $j_{1}$ and $j_{2}$.
The hyperplanes of $\mathcal{A}_{3}$ capture the transition points in $\tau^{>}$ and $\tau^{<}$. More specifically for any bijection $\rho:[m]\rightarrow[m]$, within regions that satisfy $t(\rho(1))\leq\dots\leq t(\rho(m))$, each hyperplane $T_{\rho,i,j}$ captures the points where the agent might change his choice of halting before action $i$ when the highest outcome revealed is $\rho(j)$.
Finally, the hyperplanes of $\mathcal{A}_{4}$ capture transition points in $P_{A}$. More specifically, for any two actions $i_{1}\neq i_{2}\in A$, within regions that satisfy $\tau^{>}_{i_{1}}=S_{1}$ and $\tau^{>}_{i_{2}}=S_{2}$, the hyperplane $Z_{i_{1},i_{2},S_{1},S_{2}}$ captures the points where the agent might choose to switch his ordering of $i_{1}$ and $i_{2}$.
Proposition 3.5.
[…]
The proof of this proposition relies on an analysis of each of the tuple’s components, together with Observation 3.3, and is deferred to Appendix B.5.
To conclude the proof we note that $|\mathcal{A}_{1}|=m$, $|\mathcal{A}_{2}|\leq m^{2}$, $\mathcal{A}_{3}\leq m!\cdot n\cdot m$, and $|\mathcal{A}_{4}|\leq n^{2}\cdot 2^{2m}$, so overall $|\mathcal{A}|$ is polynomial in $n$ as needed. It is also clear that the hyperplanes of $\mathcal{A}$ can be computed in polynomial time. This, together with Proposition 3.5 and Corollary 3.1 concludes the proof of Theorem 3.2. ∎
3.4 Linear vs. General Contracts
In this section we study the gap between the principal’s utility in a general contract vs. linear contract. We quantify this gap by the worst-case ratio (over all instances) between the principal’s utility in the optimal general and linear contracts.
Our main result in this section is a lower bound of $\Omega(n)$ on this gap, even for $m=3$ outcomes:
Theorem 3.3.
There exists an instance with $n$ independent actions and $m=3$ outcomes such that the worst-case ratio between the optimal principal’s utility in a general contract and in a linear contract is $\Omega(n)$.
Proof.
Consider an instance with an action set $A=[n]$, where action $i\in A$ has cost $c_{i}=\frac{2^{i}-i}{2^{n+1}}$. The outcomes are $[m]=\{1,2,3\}$, with $r(1)=0$ and $r(2),r(3)=1$. The probabilities are given by
$p_{ij}=\begin{cases}1-2^{i-n-1}&\text{if }j=1\\ c_{i}&\text{if }j=2\\ 2^{i-n-1}-c_{i}&\text{if }j=3.\end{cases}$
We first establish an upper bound on the principal’s utility from a linear contract. Fix some $\alpha\in[0,1]$, and let $t_{\alpha}$ be the corresponding linear contract.
Let $i\in A$ be some action, and consider the reservation value $z_{i}(\alpha)$, when it is non-negative (the case where it is negative is irrelevant since it implies the agent never takes action $i$). According to Equation (1) we have
$(p_{i2}+p_{i3})(\alpha-z_{i}(\alpha))=\mathbb{E}[(\alpha\cdot r(X_{i})-z_{i}(\alpha))^{+}]=c_{i}$
which implies $z_{i}(\alpha)=\alpha-\frac{c_{i}}{p_{i2}+p_{i3}}$. We now that any agent’s optimal strategy is to take actions in non-decreasing order of $\frac{c_{i}}{p_{i2}+p_{i3}}=\frac{(2^{i}-i)/2^{n+1}}{2^{i-n-1}}=1-i\cdot 2^{-i}$, i.e., in order $1,\dots,n$, until one of them reveals a non-zero outcome or $\frac{c_{i}}{p_{i2}+p_{i3}}>\alpha$. In fact, this defines a unique optimal strategy. Let $\hat{i}=\max_{i\in A:1-i\cdot 2^{-i}\leq\alpha}i$, i.e., the last action the agent is willing to take. The principal’s utility is given by
$\begin{split}u_{P}(t_{\alpha})&=(1-\alpha)\left(1-\prod_{i=1}^{\hat{i}}p_{i0}\right)=(1-\alpha)\left(1-\prod_{i=1}^{\hat{i}}\left(1-2^{i-n-1}\right)\right)\\ &\leq(1-\alpha)\left(1-\left(1-\sum_{i=1}^{\hat{i}}2^{i-n-1}\right)\right)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ (1-\alpha)2^{\hat{i}-n}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ (\hat{i}\cdot 2^{-\hat{i}})2^{\hat{i}-n}\leavevmode\nobreak\ =\leavevmode\nobreak\ \hat{i}2^{-n}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{n}{2^{n}},\end{split}$
where we’ve used that $1-\alpha\leq\hat{i}\cdot 2^{-\hat{i}}$ by the definition of $\hat{i}$.
We next establish a lower bound on the principal’s utility from the optimal general contract. Fix some $\varepsilon\in(0,1)$, and let $t$ be the contract defined by
$t(j)=\begin{cases}0&\text{if }j=1\\ 1&\text{if }j=2\\ \varepsilon&\text{if }j=3.\end{cases}$
Now, for any action $i\in A$, the reservation value $z_{i}(t)$ is given by
$\begin{split}c_{i}&=\mathbb{E}[(t(X_{i})-z_{i}(t_{\varepsilon}))^{+}]=p_{i2}(t(2)-z_{i}(t))^{+}+p_{i3}(t(3)-z_{i}(t))^{+}\\ &=c_{i}(1-z_{i}(t))^{+}+p_{i3}(\varepsilon-z_{i}(t))^{+}.\end{split}$
The only way to satisfy this equation is with $z_{i}(t)\in(0,\varepsilon)$, which gives us the equality
$\begin{split}c_{i}&=c_{i}(1-z_{i}(t))+p_{i3}(\varepsilon-z_{i}(t))=c_{i}(1-z_{i}(t))+(2^{i-n-1}-c_{i})(\varepsilon-z_{i}(t))\\ &=c_{i}+(2^{i-n-1}-c_{i})\varepsilon-2^{i-n-1}z_{i}(t).\end{split}$
This simplifies to
$z_{i}(t_{\varepsilon})=\varepsilon\left(1-\frac{c_{i}}{2^{i-n-1}}\right)=\varepsilon\cdot I\cdot 2^{-i}.$
According to Weitzman (1978), the agent takes actions in a non-increasing order of $z_{i}(t)$, i.e., in order $1,\dots,n$, and halts when one of them reveals a non-zero outcome. This implies that the principal’s utility is given by
$\begin{split}u_{P}(t)&=\sum_{i=1}^{n}\left(\prod_{k=1}^{i-1}p_{k1}\right)(p_{i2}(1-t(2))+p_{i3}(1-t(3))))=(1-t(3))\sum_{i=1}^{n}\left(\prod_{k=1}^{i-1}\left(1-2^{k-n-1}\right)\right)p_{i3}\\ &\geq(1-\varepsilon)\sum_{i=1}^{n}\left(1-\sum_{k=1}^{i-1}2^{k-n-1}\right)p_{i3}\geq(1-\varepsilon)\sum_{i=1}^{n}(1-2^{i-n-1})(2^{i-n-1}-c_{i})\\ &=(1-\varepsilon)\sum_{i=1}^{n}(1-2^{i-n-1})\frac{i}{2^{n+1}}\geq(1-\varepsilon)\frac{1}{2^{n+1}}\sum_{i=1}^{n}\frac{1}{2}i=\Omega\left(\frac{n^{2}}{2^{n}}\right).\end{split}$
To conclude, the principal’s utility from any linear contract is upper bounded by $O\left(\frac{n}{2^{n}}\right)$, while the utility from a general contract is lower bounded by $\Omega\left(\frac{n^{2}}{2^{n}}\right)$. This gives a gap of $\Omega(n)$, as desired. ∎
An upper bound.
We note that this gap is upper bounded by $O(n^{2}m)$. This upper bound is based on a result by Dütting et al. (2019), showing that the ratio between the optimal principal’s utility in a general contract and in a linear contract is upper bounded by the number of critical values of $\alpha$ in the corresponding linear contract. Combined with our Proposition 3.3, this implies an upper bound of $O(n^{2}m)$ on this ratio.
4 Correlated Actions
In this section we establish a hardness result for sequential contract design in the correlated action model. The complexity of algorithms in this model is given with respect to the size of the support of the multi-dimensional distribution from which the random variables $\{X_{i}\}_{i\in A}$ are drawn.
Our main result in this section is that in the correlated action model, it is NP-hard to approximate the optimal contract within any constant, even in the binary-outcome model (see Theorem 4.1). This result also implies a hardness result for the optimal linear contract in this setting, since in a binary-outcome model, there exists an optimal contract that is linear.
The remainder of this section is structured as follows. In Section 4.1 we formally define the binary-outcome model and present necessary preliminaries. In particular, we show an interesting equivalence between coverage set functions (see Definition 4.1) and “correlated OR” set functions (see Proposition 4.2). In Section 4.2 we present our hardness result for correlated-action sequential contracts.
4.1 Correlated Actions: Preliminaries
We restrict attention to the binary-outcome scenario. Since our result is a hardness one, this only strengthens it. In the binary-outcome scenario, the set of outcomes is $\{1,2\}$, and the corresponding rewards are $r(1)=0$ and $r(2)=1$.
In the Pandora’s Box problem corresponding to the agent’s best response problem, this means that all random variables $\{V_{i}\}_{i\in B}$ attain values in $\{0,r\}$, for some $r\in(0,1]$. We refer to this problem as the binary-outcome correlated Pandora’s box problem.
The following proposition by Chawla et al. (2020) shows that an optimal strategy in the binary-outcome correlated Pandora’s Box problem takes a simple form:
Proposition 4.1.
(Chawla et al., 2020) In the binary-outcome correlated Pandora’s Box problem any optimal strategy can be represented as a tuple $\pi=(i_{1},\dots,i_{\ell})\in B^{\ell}$ (for some $\ell\geq 0$). This tuple is interpreted as opening boxes $i_{1},\dots,i_{\ell}$ in this order, halting when a non-zero value is revealed or when all boxes in $\pi$ have been opened.
We assume without loss of generality that in such representations the last box is opened with a non-zero probability. Proposition 4.1 implies that, in a binary-outcome correlated-action model, the agent’s and principal’s utilities can be fully described by the set of actions $A$, the costs $\{c_{i}\}_{i\in A}$, and the “correlated OR” set function $f:2^{A}\rightarrow[0,1]$, where
$f(S)=Pr[\exists i\in S.\leavevmode\nobreak\ X_{i}>0].$ (2)
Indeed, the agent’s expected cost from a strategy $\pi=(i_{1},\ldots,i_{\ell})$ is given by
$c(\pi)=\sum_{k=1}^{\ell}(1-f(\{i_{1},\dots,i_{k-1}\}))c_{k}.$
Then, given a contract $t_{\alpha}$ for some $\alpha\in[0,1]$ (in the binary case, linear contracts are without loss of generality), the agent’s and principal’s expected utilities are
$u_{A}(t_{\alpha},\pi)=\alpha\cdot f(\{i_{1},\dots,i_{\ell}\})-c(\pi)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ u_{P}(t_{\alpha},\pi)=(1-\alpha)f(\{i_{1},\dots,i_{\ell}\}).$
This implies that a problem instance of the correlated-action sequential contracts can be represented by a tuple $(A,\{c_{i}\}_{i\in A},f)$. The “correlated OR” function $f$ is given by the OR of correlated Bernoulli random variables.
Definition 4.1.
A set function $f:2^{A}\rightarrow\mathbb{R}_{\geq 0}$ is coverage if there is a set of elements $U$, with associated non-negative weights $\{w_{u}\}_{u\in U}$, and a mapping $h:A\rightarrow 2^{U}$ such that for every $S\subseteq A$, $f(S)=\sum_{u\in U}w_{u}\cdot\mathbb{1}[\exists i\in S.\leavevmode\nobreak\ u\in h(i)]$. A coverage function can be described by a tuple $(U,\{w_{u}\}_{u\in U},A,h)$.
The following proposition shows an equivalence between correlated OR functions (as in Equation (2)) and coverage functions. Importantly, it also shows that the coverage function $f$ has the same (asymptotic) representation size as that of the joint distribution of $\{X_{i}\}_{i\in A}$ when given explicitly (which we defined as the size of the support). This means our choice of representation doesn’t affect an algorithm’s complexity as a function of input size.
Proposition 4.2.
For every (correlated) Bernoulli random variables $\{X_{i}\}_{i\in A}$ with support $\mathcal{R}\subseteq\{0,1\}^{A}$ and probability density function $p:\mathcal{R}\rightarrow[0,1]$, there exists a coverage function $f=(U,\{w_{u}\}_{u\in U},A,h)$ with $|U|\leq|\mathcal{R}|$ such that $f(S)=Pr[\exists i\in S.\leavevmode\nobreak\ X_{i}>0]$.
Moreover, for every coverage function $f=(U,\{w_{u}\}_{u\in U},A,h)$ such that $Im(f)\subseteq[0,1]$, there exist (correlated) Bernoulli random variables $\{X_{i}\}_{i\in A}$ with support $\mathcal{R}\subseteq\{0,1\}^{A}$ for which $f(S)=Pr[\exists i\in S.\leavevmode\nobreak\ X_{i}>0]$, and $|\mathcal{R}|\leq|U|+1$.
Proof.
Let $\{X_{i}\}_{i\in A}$ be (correlated) Bernoulli random variables with support $\mathcal{R}$ and probability density function $p:\mathcal{R}\rightarrow[0,1]$. We consider the coverage function $f=(U,\{w_{u}\}_{u\in U},A,h)$, where we define $U=\mathcal{R}$, for each $u\in U$ we define $w_{u}=p(u)$, and for each $i\in A$ we define $h(i)=\{u\in U\mid u_{i}>0\}$. This yields
$Pr[\exists i\in S.\leavevmode\nobreak\ X_{i}>0]=\sum_{v\in\mathcal{R}}p(v)\mathbb{1}[\exists i\in S.\leavevmode\nobreak\ v_{i}>0]=\sum_{u\in U}w_{u}\cdot\mathbb{1}[\exists i\in S.\leavevmode\nobreak\ u\in h(i)]=f(S),$
as needed.
In the other direction, let $f=(U,\{w_{u}\}_{u\in U},A,h)$ such that $Im(f)\subseteq[0,1]$. Assume without loss of generality [^5] This is without loss of generality because the only relevant elements of $U$ are those that belong to some $h(i)$, and the sum of their weights is $f(A)\leq 1$. that $\sum_{u\in U}w_{u}\leq 1$.
We define our variables $\{X_{i}\}_{i\in A}$ by sampling an element $x\in U\cup\{0\}$, where each element $u\in U$ is sampled with probability $Pr[x=u]=w_{u}$, and 00 is sampled with probability $1-\sum_{u\in U}w_{u}$. For any $i\in A$ and $u\in U$, we denote $X_{i}(u)=\mathbb{1}[u\in h(i)]$ and $X_{i}(0)=0$. Our random variables $\{X_{i}\}_{i\in A}$ are defined as $X_{i}=X_{i}(x)$, for any $i\in A$. Now:
$f(S)=\sum_{u\in U}w_{u}\cdot\mathbb{1}[\exists i\in S.\leavevmode\nobreak\ u\in h(i)]=\sum_{u\in U}Pr_{x}[x=u]\cdot\mathbb{1}[\exists i\in S.\leavevmode\nobreak\ X_{i}(u)>0]=Pr[\exists i\in S.\leavevmode\nobreak\ X_{i}>0],$
as needed. To show the bound on the size of $\mathcal{R}$, we note that each $X_{i}$ is a function of $x$, and there are at most $|U|+1$ possible values of $x$ we get that $\mathcal{R}$, which is defined as support of $\{X_{i}\}_{i\in A}$, has size $|\mathcal{R}|\leq|U|+1$. ∎
Remark 4.1.
Proposition 4.2 can be generalized further to an equivalence between coverage functions and “correlated maximum” functions, i.e., functions which map subsets of random variables to their expected maximum. This generalization is shown in Appendix D.
It is known that any coverage function is also submodular. A set function $f:2^{A}\rightarrow\mathbb{R}_{\geq 0}$ is submodular if for any two sets $S,S^{\prime}\subseteq A$ s.t. $S\subseteq S^{\prime}$ and $i\in A$ it holds that $f(i\mid S)\geq f(i\mid S^{\prime})$, where $f(i\mid S)=f(S\cup\{i\})-f(S)$ is the marginal contribution of $i$ to $S$.
It is also known that any submodular function (and therefore, any coverage function) is also subadditive, which means for any $S,T\subseteq A$, $f(S\cup T)\leq f(S)+f(T)$.
4.2 Hardness of Approximation of the Optimal Contract Under Correlated Actions
Chawla et al. (2020) established a hardness of approximation for the correlated Pandora’s Box problem. However, this hardness result does not directly imply a hardness of approximation for the optimal contract problem. This is because the hardness of Pandora’s Box corresponds to the hardness of the agent’s best response problem, but even if the agent’s problem is hard to approximate, it may still be the case that approximating the optimal contract admits a poly-time algorithm.
In this section we present the following hardness of approximation result.
Theorem 4.1.
In the correlated action model, it is NP-hard to approximate the optimal contract within any constant, even in the binary-outcome case.
The hardness we reduce from is an NP-hard promise problem dealing with coverage functions, introduced in (Ezra et al., 2024). This promise problem is described in the following proposition:
Proposition 4.3 ((Ezra et al., 2024)).
For every $M>1$ and every $0<\varepsilon<e^{-1}$, on input $(k,f)$, where $k\in\mathbb{N}$ and $f=(U,\{w_{u}\}_{u\in U},A,h)$ is a coverage function with with $Im(f)\subseteq[0,1]$, such that $\forall i\in A.\leavevmode\nobreak\ f(\{i\})=\frac{1}{k}$ and exactly one of the following two conditions holds:
  1. 1.
    There exists a set $S\subseteq A$ of size $k$ such that $f(S)=1$.
  2. 2.
    Every set $S\subseteq A$ of size $\beta k$ such that $\beta\leq M$ satisfies $f(S)\leq 1-e^{-\beta}+\varepsilon$.
It is NP-hard to determine which of the two conditions is satisfied by the input.
To prove Theorem 4.1, for any constant $\gamma\in(0,1)$, we show a reduction from the promise problem in Proposition 4.3 such that, for any $k>3$, under case (1) of Proposition 4.3 the optimal contract has utility at least $\frac{1}{4}$, and under case (2) any contract $t_{\alpha}$ with $\alpha<1-\frac{\gamma}{8}$ has a utility of 00, but the contract $\alpha=1-\frac{\gamma}{8}$ has a positive utility. This implies that, when $t_{\alpha}$ is a $\gamma$-approximation of the optimal contract, under the first case it holds that $\alpha\leq 1-\frac{\gamma}{4}$ (since the principal’s utility is upper-bounded by $1-\alpha$, and $u_{P}(t_{\alpha})\geq\gamma\cdot\frac{1}{4}$), while under the second it holds that $\alpha\geq 1-\frac{\gamma}{8}$, which rules out a polytime $\gamma$-approximation algorithm.
Proof of Theorem 4.1.
For any $(k,f^{\prime}:2^{A^{\prime}}\rightarrow[0,1])$ per Proposition 4.3 with the parameters $M=3,\varepsilon=0.001$, where $f^{\prime}=(U,\{w_{u}\}_{u\in U},A^{\prime},h^{\prime})$. Construct the following instance of the sequential contracts problem: $(A,\{c_{i}\}_{i\in A},f)$ defined as follows. Our set of actions is $A=A^{\prime}\cup\{0\}$. Our costs are $c_{i}=\frac{1.5}{k+1}$ for any $i\in A^{\prime}$ and $c_{0}=1-\frac{\gamma}{8}$. Our “correlated OR” function is $f=(U,\{w_{u}\}_{u\in U},A,h)$, where $h$ is defined as
$h(i)=\begin{cases}h^{\prime}(i)&i\neq 0\\ U&i=0.\end{cases}$
Note that for any $S\subseteq A$ that doesn’t contain 00 we have $f(S)=f^{\prime}(S)$, while for any $S\subseteq A$ that contains 00 we have $f(S)=1$.
We first note the following Lemma:
Lemma 4.1.
For any $\alpha<c_{0}$, the agent’s best response to $t_{\alpha}$ never contains 00.
Proof.
Let $\pi=(s_{1},\dots,s_{\ell})$ be the agent’s best response to the contract $t_{\alpha}$, and assume by contradiction that it contains action 00. It is clear that 00 is the final action, due to our assumption that in the strategy representation the final action must be taken with non-zero probability.
Denote $\pi^{\prime}=(s_{1},\dots,s_{\ell-1})$, we show that $\pi^{\prime}$ has a greater utility than $\pi$, leading to a contradiction. Indeed,
$\begin{split}c(\pi)-c(\pi^{\prime})&=\sum_{i=1}^{\ell}(1-f(\{s_{1},\dots,s_{i-1}\}))\cdot c_{s_{i}}-\sum_{i=1}^{\ell-1}(1-f(\{s_{1},\dots,s_{i-1}))\cdot c_{s_{i}}\\ &=(1-f(\{s_{1},\dots,s_{\ell-1}\}))\cdot c_{0},\end{split}$
which means
$\begin{split}u_{A}(t_{\alpha},\pi)-u_{A}(t_{\alpha},\pi^{\prime})&=\alpha\cdot f(\{s_{1},\dots,s_{\ell-1}\})-c(\pi)-(\alpha\cdot f(\{s_{1},\dots,s_{\ell}\})-c(\pi^{\prime}))\\ &=(c_{0}-\alpha)\cdot(1-f(\{s_{1},\dots,s_{\ell-1}\}))<0,\end{split}$
where the last inequality is because the probability of the last action (i.e., $s_{\ell}$) being taken is non-zero. This contradicts $\pi$ being a best response. ∎
We now prove the correctness of our reduction.
Case 1: $f^{\prime}$ satisfies condition (1) from Proposition 4.3.
We bound the principal’s utility from the optimal contract from below by her utility from the contract $t_{\frac{3}{4}}$.
Let $S$ be a set that satisfies condition $1$ (i.e., $f^{\prime}(S)=1$ and $|S|=k$). Note that for any set $T\subseteq S$ it holds that $f^{\prime}(T)=\frac{|T|}{k}$. The upper bound is from subadditivity, and since for any $i\in A^{\prime}$ it holds that $f^{\prime}(\{i\})=\frac{1}{k}$, and the lower bound is since, from subadditivity, $f^{\prime}(T)\geq f^{\prime}(S)-f^{\prime}(S\setminus T)\geq 1-\frac{|S|-|T|}{k}$. Now, let $s_{1},\dots,s_{k}\in A$ be some enumeration of the elements of $S$, and consider the strategy $\pi=(s_{1},\dots,s_{k})$, we claim that this maximizes the agent’s utility. The expected cost $c(\pi)$ is
$\begin{split}c(\pi)&=\sum_{i=1}^{k}(1-f(\{s_{1},\dots,s_{i-1}\}))c_{s_{i}}=\sum_{i=1}^{k}\left(1-\frac{i-1}{k}\right)\frac{1.5}{k+1}=\frac{1.5}{k+1}\left(k-\frac{1}{k}\sum_{i=1}^{k}(i-1)\right)\\ &=\frac{1.5}{k+1}\left(k-\frac{1}{k}\frac{k(k-1)}{2}\right)=\frac{1.5}{k+1}\frac{2k-(k-1)}{2}=\frac{3}{4}.\end{split}$
This means that the agent’s utility from $\pi$ is
$u_{A}(t_{\frac{3}{4}},\pi)=\frac{3}{4}f(S)-c(\pi)=0.$
Now, let $\pi^{\prime}=(a_{1},\dots,a_{\ell})$ be an agent’s strategy which doesn’t contain 0 (this is without loss of generality due to Lemma 4.1 and because $\frac{3}{4}<1-\frac{\gamma}{8}$). We show that the agent’s utility from $\pi^{\prime}$ is non-positive, which, due to tie-breaking in favor of the principal, means that $\pi$ (or some other strategy which leads to a success with a probability of 1) is the agent’s best response.
If $\ell\leq k$, then
$\begin{split}c(\pi^{\prime})&=\sum_{i=1}^{\ell}(1-f(\{a_{1},\dots,a_{i-1}\}))c_{a_{i}}\geq\sum_{i=1}^{\ell}\left(1-\frac{i-1}{k}\right)\frac{1.5}{k+1}=\frac{1.5}{k+1}\left(\ell-\frac{1}{k}\sum_{i=1}^{\ell}(i-1)\right)\\ &=\frac{1.5}{k+1}\left(\ell-\frac{1}{k}\frac{\ell(\ell-1)}{2}\right)=\frac{1.5}{k+1}\left(\frac{\ell(2k-(\ell-1))}{2k}\right)\geq\frac{1.5\cdot\ell}{2k},\end{split}$
where the first inequality is due to subadditivity, for any $S\subseteq A^{\prime}$ it holds that $f(S)=f^{\prime}(S)\leq\frac{|S|}{k}$, and the last inequality is because $\ell\leq k$. This means the agent’s utility from $\pi^{\prime}$ is at most
$u_{A}\left(t_{\frac{3}{4}},\pi^{\prime}\right)=\frac{3}{4}f(\{a_{1},\dots,a_{\ell}\})-c(\pi^{\prime})\leq\frac{3}{4}\frac{\ell}{k}-\frac{3}{4}\frac{\ell}{k}=0,$
as needed.
If $\ell>k$, we can bound $c(\pi^{\prime})$ from below by $c((a_{1},\dots,a_{k}))$, which as argued above is at least $\frac{3}{4}$, which means the agent’s utility from $\pi^{\prime}$ is at most
$u_{A}\left(t_{\frac{3}{4}},\pi^{\prime}\right)=\frac{3}{4}f(\{a_{1},\dots,a_{k}\})-c(\pi^{\prime})\leq\frac{3}{4}\cdot 1-\frac{3}{4}=0,$
as needed.
Because $\pi$ maximizes the agent’s best response, and because the agent breaks ties in favor of the principal, the principal’s utility from $t_{\frac{3}{4}}$ is at least
$u_{P}\left(t_{\frac{3}{4}}\right)\geq\left(1-\frac{3}{4}\right)f(S)=\frac{1}{4},$
as needed.
Case 2: $f$ satisfies condition (2) from Proposition 4.3.
It is easy to see that the agent’s best response to a contract $t_{1-\frac{\gamma}{8}}$ is $(0)$, which means $u_{P}(t_{1-\frac{\gamma}{8}})=\frac{\gamma}{8}>0$. We conclude the proof by showing that under a contract $t_{\alpha}$ with $\alpha<1-\frac{\gamma}{8}$, the agent’s utility from any non-empty strategy which doesn’t include 00 is negative. This means that the agent’s best response to $t_{\alpha}$ is to do nothing, implying $u_{P}(t_{\alpha})=0$, as needed.
Let $\alpha<1-\frac{\gamma}{8}$, and consider some strategy $(s_{1},\dots,s_{\ell})$ which doesn’t include 00.
If $\ell<\frac{k}{10}$, the expected cost of strategy $\pi$ is at least
$c(\pi)\geq\frac{1.5}{k+1}\left(\frac{\ell(2k-(\ell-1))}{2k}\right)>\frac{1.5\cdot\ell}{2k}\left(\frac{(2-\frac{1}{10})k}{k+1}\right),$
where the first inequality is the same as we’ve seen in the analysis of case 1. For $k\geq 3$ this implies $c(\pi)>\frac{\ell}{k}$, which means the agent’s utility from $\pi$ is bounded by
$u_{A}(t_{\alpha},\pi)=\alpha\cdot f(\{s_{1},\dots,s_{\ell}\})-c(\pi)<1\cdot\frac{\ell}{k}-\frac{\ell}{k}<0,$
where the inequality is from our earlier observation that $f(T)\leq\frac{|T|}{k}$, and from our bound on $c(\pi)$.
If $\frac{k}{10}\leq\ell\leq 3k$, by applying condition (2) we can bound the expected cost $c(\pi)$ from below by
$\begin{split}c(\pi)&=\sum_{i=1}^{\ell}(1-f(\{s_{1},\dots,s_{i-1}\})c_{s_{i}}\geq\sum_{i=1}^{\ell}(e^{-(i-1)/k}-\varepsilon)c_{s_{i}}=\frac{1.5}{k+1}\left(\sum_{i=1}^{\ell}e^{-(i-1)/k}-\ell\varepsilon\right)\\ &=\frac{1.5}{k+1}\left(\frac{1-e^{-\ell/k}}{1-e^{-1/k}}-\ell\varepsilon\right)>1.5\left(\frac{1-e^{-\ell/k}}{(1-e^{-1/k})(k+1)}-3\varepsilon\right)>1.5\left(\frac{1-e^{-\ell/k}}{1.3}-3\varepsilon\right)\\ &>1-e^{-\ell/k}+\varepsilon,\end{split}$
where the second to last inequality is because $(1-e^{-x})(x+1)<1.3$ for any $x>0$, and the last inequality is because our choice of $\varepsilon$ is such that for $\ell\geq\frac{k}{10}$ the inequality is satisfied. This implies
$u_{A}(t_{\alpha},\pi)=\alpha\cdot f(\{s_{1},\dots,s_{\ell}\})-c(\pi)<\alpha(1-e^{-\ell/k}+\varepsilon)-(1-e^{-\ell/k}+\varepsilon)<0.$
If $\ell>3k$, we can bound the cost of $c(\pi)$ from below by the cost of $c((s_{1},\dots,s_{3k}))$, which, as argued above is at least
$1.5\left(\frac{1-e^{-3}}{1.3}-3\varepsilon\right)>1,$
implying the agent’s utility is negative. This concludes the proof of Theorem 4.1. ∎
Appendix A Polynomial Time Algorithm For the Principal’s Utility Under Independent Actions
In this appendix we prove Proposition 3.2. See 3.2
The first challenge we address is that while finding an agent’s optimal strategy is easy, it is unclear how to efficiently find an agent’s optimal strategy that leads to the best principal’s utility. To overcome this, for a given contract $t$, we devise a contract $t^{\varepsilon}$ which is sufficiently close to $t$ such that maximizing the agent’s utility under $t^{\varepsilon}$ also maximizes his utility under $t$, but also causes the agent’s utility itself to break ties in favor of the principal. This contract is established in the following proposition, and we later analyze its properties.
Proposition A.1.
For a small enough $\varepsilon>0$, it holds that $t^{\varepsilon}(j)=t(j)+\varepsilon\cdot(r(j)-t(j))$ is a valid contract (i.e., all payments are non-negative), and any strategy $\pi$ which maximizes the agent’s utility under $t^{\varepsilon}$ also maximizes his utility under $t$.
Proof.
It is easy to see that $t^{\varepsilon}$ is a valid contract for a small enough $\varepsilon>0$, Indeed, if $t(j)$ is positive, for a small enough $\varepsilon$, it holds that $t^{\varepsilon}(j)$ is non-negative. On the other hand, if $t(j)=0$, then $t^{\varepsilon}(j)\geq t(j)\geq 0$.
From Observation 3.1, it suffices to show that for a small enough $\varepsilon$ the following conditions are satisfied:
  1. 1.
    For any $i_{1},i_{2}$ such that $z_{i_{1}}(t)<z_{i_{2}}(t)$ it holds that $z_{i_{1}}(t^{\varepsilon})<z_{i_{2}}(t^{\varepsilon})$.
  2. 2.
    For any $i,j$ such that $t(j)>z_{i}(t)$ it holds that $t^{\varepsilon}(j)>z_{i}(t^{\varepsilon})$.
  3. 3.
    For any $i,j$ such that $t(j)<z_{i}(t)$ it holds that $t^{\varepsilon}(j)<z_{i}(t^{\varepsilon})$.
  4. 4.
    For any $j_{1},j_{2}$ such that $t(j_{1})<t(j_{2})$ it holds that $t^{\varepsilon}(j_{1})<t^{\varepsilon}(j_{2})$.
From the continuity of $z_{i}(t)$, it is immediate that these conditions hold for a small enough $\varepsilon$. ∎
Furthermore, a small enough $\varepsilon$ that satisfies Proposition A.1 can be found. The following lemma bounds the change in $z_{i}(t)$ as $t$ increases/decreases slightly, which directly implies how to find an $\varepsilon$ that satisfies the inequalities described in the proof of Proposition A.1.
Lemma A.1.
For any two contracts $t,t^{\prime}$ and any $\delta>0$, if for any outcome $j\in[m]$ it holds that
$t(j)-\delta\leq t^{\prime}(j)\leq t(j)+\delta,$
then it also holds that for any action $i\in A$
$z_{i}(t)-\delta\leq z_{i}(t^{\prime})\leq z_{i}(t)+\delta.$
Proof.
Let $i\in A$ be some action. By the definition of the reservation value (see Equation (1)), $z_{i}(t^{\prime})$ is such that satisfies
$\mathbb{E}[(t^{\prime}(X_{i})-z_{i}(t^{\prime}))^{+}]=c_{i},$
and $z_{i}(t)$ satisfies
$\mathbb{E}[(t(X_{i})-z_{i}(t))^{+}]=c_{i}.$
We start with the lower bound. To show that $z_{i}(t^{\prime})\geq z_{i}(t)-\delta$ it suffices to show that the inequality $\mathbb{E}[(t^{\prime}(X_{i})-(z_{i}(t)-\delta))^{+}]\geq c_{i}$ holds. Indeed,
$\mathbb{E}[(t^{\prime}(X_{i})-(z_{i}(t)-\delta))^{+}]\geq\mathbb{E}[(t(X_{i})-\delta-z_{i}(t)+\delta)^{+}]=c_{i}.$
For the upper bound, we show that $\mathbb{E}[(t^{\prime}(X_{i})-(z_{i}(t)+\delta))^{+}]\leq c_{i}$. Similarly to before we have
$\mathbb{E}[(t^{\prime}(X_{i})-(z_{i}(t)+\delta))^{+}]\leq\mathbb{E}[(t(X_{i})+\delta-z_{i}(t)-\delta)^{+}]=c_{i},$
concluding the proof. ∎
The following lemma shows an important relation between the agent’s utilities in $t$ and $t^{\varepsilon}$, and the principal’s utility in $t$. This relation is the reason we examine $t^{\varepsilon}$.
Lemma A.2.
For any contract $t$, any $\varepsilon>0$ such that $t^{\varepsilon}$ is a valid contract, and any agent’s strategy $\pi$ it holds that
$u_{A}(t^{\varepsilon},\pi)=\varepsilon\cdot u_{P}(t,\pi)+u_{A}(t,\pi).$
Proof.
Recall that $S(\pi)$ denotes the random variable depicting the set of actions taken by the agent under $\pi$, and $\omega(\pi)$ is the random variable depicting the project’s final outcome under $\pi$ (see Section 2).
$\begin{split}u_{A}(t^{\varepsilon},\pi)&=\mathbb{E}[t^{\varepsilon}(\omega(\pi))-\sum_{i\in S(\pi)}c_{i}]=\mathbb{E}[\varepsilon\cdot(r(\omega(\pi))-t(\omega(\pi)))]+\mathbb{E}[t(\omega(\pi))-\sum_{i\in S(\pi)}c_{i}]\\ &=\varepsilon\cdot u_{P}(t,\pi)+u_{A}(t,\pi),\end{split}$
as needed. ∎
As a corollary of Proposition A.1 and Lemma A.2 we get
Corollary A.1.
For any contract $t$, let $t^{\varepsilon}$ be per Proposition A.1. Any agent’s optimal strategy $\pi$ under the contract $t^{\varepsilon}$ satisfies $u_{P}(t,\pi)=u_{P}(t)$.
To address the challenge of computing the principal’s utility under some strategy, we consider a simple type of agent’s strategies for which we can efficiently compute the principal’s utility.
Definition A.1.
An agent’s strategy $\pi$ is non-adaptive if it can be described by a tuple $(\sigma,\rho,\tau)$, where
  1. 1.
    $\sigma:[n]\rightarrow A$ is a bijection which defines the action taking order in the following way: if $\pi$ doesn’t halt before it, the $i$th action is taken is $\sigma(i)$.
  2. 2.
    $\rho:[m]\rightarrow[m]$ is a bijection describing the agent’s preference of outcomes, i.e., the agent prefers outcome $j_{1}$ over $j_{2}$ if $\rho(j_{1})>\rho(j_{2})$.
  3. 3.
    $\tau=\{\tau_{i}\}_{i\in A}\subseteq[m]$ defines the halting condition in the following way: before taking an action $i\in A$, $\pi$ halts if his current preferred outcome $j$ satisfies $\rho(j)\geq\rho(\tau_{i})$.
  4. 4.
    Upon halting $\pi$ chooses its preferred outcome according to $\rho$ as the project’s final outcome.
By Observation 3.1 it is easy to see that a non-adaptive optimal strategy exists, and can be easily found [^6] Let $\sigma$ describe some ordering of the actions such that the reservation value is non-increasing, Let $\rho$ be some ordering that is non-increasing, and let $\tau_{i}\in\text{argmax}_{j\in[m]:t(j)<z_{i}(t)}t(j)$..
For any non-adaptive strategy $\pi=(\sigma,\tau,\mu)$ one can compute the probability of the final outcome being $j$ for any $j\in[m]$ by iterating over the actions taken, according to the order $\sigma$, while maintaining the probability of $j$ being the agent’s preferred revealed outcome so far. From this distribution over the outcomes, the principal’s utility is immediate.
Appendix B Additional Proofs from Section 3
In this appendix we provide technical details on proofs from Section 3.
B.1 Proof of Lemma 3.1
In this section we present the proof of Lemma 3.1, See 3.1
Proof.
If $c_{i}=0$, $z_{i}(\alpha)=\infty$, and we are done.
Otherwise, if $c_{i}>0$, for any $\alpha\in[0,1]$, the random variable $t_{\alpha}(X_{i})$ attains values within $\{\alpha\cdot r(j)\}_{j\in[m]}$, and equals $\alpha\cdot r(j)$ with probability $p_{ij}$. Thus, $z_{i}(\alpha)$ is given by
$\sum_{j\in[m]:\alpha r(j)>z_{i}(\alpha)}p_{ij}(\alpha\cdot r(j)-z_{i}(\alpha))=c_{i}.$ (3)
Now, for any $j\in[m]$ we define $\alpha_{i,j}=\inf\{\alpha\in[0,1]\mid z_{i}(\alpha)\geq\alpha\cdot r(j)\}$, which we interpret as $\infty$ if the set is empty. Intuitively, $\alpha_{i,j}$ is the threshold of $\alpha$ from which the agent is willing to execute action $i$ when he has already revealed outcome $j$. We also denote $\alpha_{i,0}=0$. Observe that $\alpha_{i,0}\leq\alpha_{i,1}\leq\dots\leq\alpha_{i,m}$, and that $\alpha_{i,m}=\infty$ (since $z_{i}(\alpha)$ can never be greater than $\alpha\cdot r(m)$ for $i$ such that $c_{i}>0$).
For any $j\in[m]$, such that $\alpha_{i,(j-1)}\neq\alpha_{i,j}$ we consider the interval $[\alpha_{i,(j-1)},\alpha_{i,j})$. For any $\alpha\in[\alpha_{i,(j-1)},\alpha_{i,j})$, the set $\{j\in[m]\mid\alpha\cdot r(j)>z_{i}(\alpha)\}$ is precisely $\{j,\dots,m\}$. Plugging this into Equation (3) gives us
$\sum_{k=j}^{m}p_{ik}(\alpha\cdot r(k)-z_{i}(\alpha))=c_{i},$
implying that within the interval $[\alpha_{i,(j-1)},\alpha_{i,j})$,
$z_{i}(\alpha)=\frac{\sum_{k=j}^{m}p_{ik}r(k)}{\sum_{k=j}^{m}p_{ik}}\cdot\alpha-\frac{c_{i}}{\sum_{k=j}^{m}p_{ik}}.$
Therefore, $z_{i}(\alpha)$ is piecewise linear and its slope is monotonically increasing.
We have shown at most $m$ intervals, within each of which $z_{i}(\alpha)$ is linear, and such that from one interval to the next the slope is weakly increasing, which proves our desired piecewise linearity and convexity. Since we’ve given an explicit expression for $z_{i}(\alpha)$ within each interval $[\alpha_{i,(j-1)},\alpha_{i,j})$, the only missing piece for an explicit expression for $z_{i}(\alpha)$ is explicit closed-form expressions for $\alpha_{i,j}$ for any $j\in[m]$. Indeed, an explicit expression for $\alpha_{i,j}$ follows from the continuity of $z_{i}(\alpha)$. By definition of $\alpha_{i,j}$, because $z_{i}(\alpha)$ is continuous, the following equality holds
$z_{i}(\alpha)=\alpha_{i,j}\cdot r(j),$
which, in combination with Equality 3, gives
$\alpha_{i,j}=\frac{c_{i}}{\sum_{k=j}^{m}p_{ik}(r(k)-r(j))}.$
Note that all of our closed-form expressions are efficiently computable, concluding the proof. ∎
B.2 Proof of Lemma 3.2
In this section we provide the proof for Lemma 3.2. See 3.2
Proof.
Let $\alpha^{\prime}\in(0,1)$ be some critical value. Denote $\pi=\pi^{\star}(t_{\alpha^{\prime}})$. According to Observation 3.2, $\pi$ isn’t an optimal strategy directly to the left of $\alpha^{\prime}$. Consider which condition of Observation 3.1 is violated by $\pi$ directly to the left of $\alpha^{\prime}$.
Assume condition (1) of Observation 3.1 is violated by $\pi$ to the left of $\alpha^{\prime}$, i.e., there exist actions $i_{1}\neq i_{2}\in A$ such that with non-zero probability when action $i_{1}$ is taken, $i_{2}$ is still untaken and $z_{i_{1}}(\alpha)<z_{i_{2}}(\alpha)$. Because $\pi$ is optimal at $\alpha^{\prime}$, we have $z_{i_{1}}(\alpha^{\prime})\geq z_{i_{2}}(\alpha^{\prime})$, which implies (from continuity) the first condition of Lemma 3.2 holds, as needed.
Assume condition (2) of Observation 3.1 is violated, i.e., with non-zero probability when $j\in[m]$ is a revealed outcome of maximum payment and $i\in A$ is an untaken action of maximum $z_{i}(\alpha)$, it holds that either $\alpha\cdot r(j)<z_{i}(\alpha)$ and $\pi$ halts or $\alpha\cdot r(j)>z_{i}(\alpha)$ and $\pi$ doesn’t halt. If $\alpha\cdot r(j)<z_{i}(\alpha)$ and $\pi$ halts, then because $\pi$ is optimal at $\alpha^{\prime}$, we have $\alpha^{\prime}\cdot r(j)\geq z_{i}(\alpha^{\prime})$, which implies the second condition of Lemma 3.2, as needed. Otherwise, if $\alpha\cdot r(j)>z_{i}(\alpha)$ and $\pi$ doesn’t halt, then because $\pi$ is optimal at $\alpha^{\prime}$ we have $\alpha^{\prime}\cdot r(j)\leq z_{i}(\alpha^{\prime})$, which also implies the second condition of Lemma 3.2, as needed.
Finally, assume towards contradiction that condition (3) of Observation 3.1 is violated to the left of $\alpha^{\prime}$, i.e. with non-zero probability upon halting $\alpha\cdot r(\omega(\pi))<\alpha\cdot r(j)$ for some revealed outcome $j$. This implies $r(\omega(\pi))<r(j)$, which implies $\alpha^{\prime}\cdot r(\omega(\pi))<\alpha^{\prime}\cdot r(j)$, contradicting $\pi$ being optimal at $\alpha^{\prime}$. ∎
B.3 Proof of Lemma 3.4
In this section we prove Lemma 3.4. See 3.4
Proof.
We begin by noting that the principal’s utility is bounded(for example by the agent’s utility under the linear contract $t_{1}$), therefore it has a supremum in $Q$. We denote this supremum by $u$, i.e., $u=\sup\{u_{P}(t)\mid t\in Q\}$.
Let $\{t^{\ell}\}_{\ell\in\mathbb{N}}$ be a sequence in $Q$ with utilities that converge to $u$. Since $Q$ is a polytope, and therefore bounded, the sequence $\{t^{\ell}\}_{\ell\in\mathbb{N}}$ has a converging subsequence, assume without loss of generality that it is $\{t^{\ell}\}_{\ell\in\mathbb{N}}$ itself, and denote its limit by $t^{\prime}$. By showing $u_{P}(t^{\prime})\geq u$, we are done.
Because the number of possible agent’s strategies is finite, there exists a strategy $\pi$ and subsequence of $\{t^{\ell}\}_{\ell\in\mathbb{N}}$ in which the agent’s best response is always $\pi$. Assume without loss of generality that this subsequence is $\{t_{\ell}\}_{\ell\in\mathbb{N}}$ itself. Since for any strategy $\pi^{\prime}$ it holds that $u_{A}(t,\pi^{\prime})$ is continuous (in $t$), and since $\pi$ maximizes the agent’s utility in every $t_{\ell}$, it also maximizes the agent’s utility in $t^{\prime}$, which due to tie-breaking in favor of the principal means
$u_{P}(t^{\prime})=u_{P}(t^{\prime},\pi^{\star}(t^{\prime}))\geq u_{P}(t^{\prime},\pi)=u,$
where the last equality is due to the continuity (in $t$) of $u_{P}(t,\pi$). ∎
B.4 Proof of Lemma 3.5
In this section we prove Lemma 3.5 See 3.5
Proof.
Let $t^{0}\in Q$ be such that maximizes the principal’s utility in $Q$. If $t^{0}\notin\text{relint}(Q)$, we are done. Otherwise, we show a contract $t^{1}$ in the relative boundary of $Q$ with utility $u_{P}(t^{1})\geq u_{P}(t^{0})$. Note that for any agent’s strategy $\pi$, the principal’s utility from a contract $t$ when the agent engages in strategy $\pi$ is given by the following linear (in $t$) expression:
$u_{P}(t,\pi)=\sum_{j\in[m]}P_{j}(r(j)-t(j)),$
where $P_{j}$ is the probability of the project’s outcome being $j$ when the agent engages in strategy $\pi$. It is known that a linear expression is maximized at a vertex of a polytope, meaning $u_{P}(t,\pi^{\star}(t^{0}))$ is maximized at some $t^{1}$ which is a vertex of $Q$ and therefore in its relative boundary. Since $\pi^{\star}(t^{0})$ maximizes the agent’s utility in $t^{0}\in\text{relint}(Q)$, due to our assumption that $(P_{A},P_{[m]},\tau^{>},\tau^{<})$ is constant in $\text{relint}(Q)$, we get that $\pi^{\star}(t^{0})$ maximizes the agent’s utility everywhere in $\text{relint}(Q)$. Because, when fixing any strategy $\pi$, the agent’s utility $u_{A}(t,\pi)$ is continuous in $t$, we get that $\pi^{\star}(t_{0})$ also maximizes the agent’s utility in $t_{1}$. Due to tie-breaking in favor of the principal, it holds that
$u_{P}(t^{1})=u_{P}(t^{1},\pi^{\star}(t^{1}))\geq u_{P}(t^{1},\pi^{\star}(t^{0}))\geq u_{P}(t^{0},\pi^{\star}(t^{0}))=u_{P}(t^{0}),$
as needed. ∎
B.5 Proof of Proposition 3.5
In this appendix we prove Proposition 3.5 See 3.5
Proof.
Let $F$ be some face of the hyperplane arrangement such that $F\subseteq[0,L]^{m}$, and let $t^{1},t^{2}\in\text{relint}(F)$.
Assume by contradiction that $P_{[m]}(t^{1})\neq P_{[m]}(t^{2})$. This means there exist some $j_{1},j_{2}\in[m]$ such that $j_{1}\leq_{P_{[m]}(t^{1})}j_{2}$ but $j_{1}\not\leq_{P_{[m]}(t^{2})}j_{2}$. Equivalently, $t^{1}(j_{1})\leq t^{1}(j_{2})$ but $t^{2}(j_{1})>t^{2}(j_{2})$. This implies the existence of $t\in\text{relint}(F)$ such that $t(j_{2})=t(j_{1})$, meaning $t$ is in $E_{j_{1},j_{2}}$ but $t^{2}$ is not, contradicting Observation 3.3.
We have that $P_{[m]}$ is constant within $\text{relint}(F)$, which means there exists some bijection $\rho:[m]\rightarrow[m]$ such that
$t(\rho(1))\leq\dots\leq t(\rho(m))$
for any $t\in\text{relint}(F)$.
Assume by contradiction that $\tau^{>}(t^{1})\neq\tau^{>}(t^{2})$. This implies the existence of $i\in A$ and $j\in[m]$ such that $t^{1}(\rho(j))\leq z_{i}(t^{1})$ but $t^{2}(\rho(j))>z_{i}(t^{2})$. Note that $t^{1}(\rho(j))\leq z_{i}(t^{1})$ implies
$\sum_{k=j}^{m}p_{i\rho(k)}(t^{1}(\rho(k))-t^{1}(\rho(j))\geq c_{i},$
and $t^{2}(\rho(j))>z_{i}(t^{2})$ implies
$\sum_{k=j}^{m}p_{i\rho(k)}(t^{2}(\rho(k))-t^{2}(\rho(j))<c_{i}.$
This implies the existence of $t\in\text{relint}(F)$ such that $t\in T_{\rho,i,j}$ but $t^{2}\notin T_{\rho,i,j}$, contradicting Observation 3.3.
The proof for $\tau^{<}$ being constant is almost entirely the same as what we did for $\tau^{>}$ above.
Finally, assume by contradiction that there exist $t^{1},t^{2}\in\text{relint}(F)$ with $P_{A}(t^{1})\neq P_{A}(t^{2})$. This implies the existence of $i_{1},i_{2}\in A$ such that $z_{i_{1}}(t^{1})\leq z_{i_{2}}(t^{1})$ but $z_{i_{1}}(t^{2})>z_{i_{2}}(t^{2})$. From continuity, there exists $t\in\text{relint}(F)$ such that $z_{i_{1}}(t)=z_{i_{2}}(t)$, which means $t\in Z_{i_{1},i_{2},\tau^{>}_{i_{1}},\tau^{>}_{i_{2}}}$ but $t^{2}\notin Z_{i_{1},i_{2},\tau^{>}_{i_{1}}}$, contradiction. ∎
Appendix C Superpolynomial Number of Best Responses
In this section, we show that there exist instances with independent actions that admit super-polynomially many best responses of the agent, when $m$ is not a constant. Specifically, we prove the following proposition:
Proposition C.1.
For any $m\geq 2$, and any $n\in\mathbb{N}$, there exists an instance of the independent-action sequential contract problem with $n$ actions and $m$ outcomes such that there are $n^{\Omega(m)}$ best responses to monotone contracts.
Proof.
Let $m\geq 2$ and let $n\in\mathbb{N}$. Assume without loss of generality [^7] Otherwise we can take the maximal $n^{\prime}<n$ such that $m-1$ divides $n^{\prime}$, and add $n-n^{\prime}$ “dummy” actions. that $m-1$ divides $n$, and let $\ell=\frac{n}{m-1}$. We construct the following instance:
  1. 1.
    The set of outcomes is $[m]$, and for any outcome $j\in\{2,\dots,m\}$, the reward is $r(j)=\ell^{j}$.
  2. 2.
    The set of actions is $A=\{2,\dots,m\}\times[\ell]$, and for every action $a=(j,i)\in A$, $c_{a}=i\cdot r(j)$.
  3. 3.
    For any action $a=(j,i)$ and any outcome $j^{\prime}\in[m]$, $p_{aj^{\prime}}$ is given by
    $p_{aj^{\prime}}=\begin{cases}\frac{1}{2}&\mbox{if }j^{\prime}=1\lor j^{\prime}=j\\ 0&\mbox{else}.\end{cases}$
Essentially, this definition allows us the following two degrees of freedom in the best response, by a careful choice of contracts. Firstly, for any outcome $\{2,\dots,m\}$, we can choose how many actions from $\{j\}\times[\ell]$ are taken with non-zero probability. Secondly, we can choose the order in which the actions $\{2,\dots,m\}\times\{1\}$ are taken. When $m$ is small, we will use the former, and when $m$ is large, the latter. We begin our case by case analysis:
Case 1: $m\leq\sqrt{n}$.
For any vector $v\in[\ell]^{m-1}$, let $t_{v}$ denote the following contract:
$t_{v}(j)=\begin{cases}0&\mbox{if }j=1\\ (2+\frac{1}{2\ell})r(j)\cdot v_{j-1}&\mbox{if }j>1.\end{cases}$
Note that for any $v\in[\ell]^{m-1}$, this contract is monotone. We claim that for every $v\in[\ell]^{m-1}$, the agent’s best response to $t_{v}$, is different. This follows from the following observation, which means that the set of actions taken with non-zero probability is different.
Observation C.1.
Let $v\in[\ell]^{m-1}$. Consider the agent’s best response to $t_{v}$. Let $a=(j,i)$ be some action, it is easy to see that if $i\leq v_{j-1}$ then $z_{a}(t_{v})>0$, meaning action $a$ is taken with non-zero probability, and if $i>v_{j-1}$ then $z_{a}(t_{v})<0$, meaning action $a$ is never taken.
This implies that the number of best responses is lower bounded by
$\ell^{m-1}\geq(\sqrt{n})^{m-1}=n^{\Omega(m)},$
as needed.
Case 2: $m>\sqrt{n}$.
For any bijection $\rho:[m-1]\rightarrow\{2,\dots,m\}$, denote by $t_{\rho}$ the following contract:
$t_{\rho}(j)=\begin{cases}0&\mbox{if }j=1\\ \rho^{-1}(j)+2r(j)&\mbox{if }j>1.\end{cases}$
Note that for any bijection $\rho$, $t_{\rho}$ is monotone. We claim that for each $\rho$ the agent’s best response to $t_{\rho}$ has a different action-taking order. This is implied by the following observation:
Observation C.2.
Let $\rho:[m-1]\rightarrow\{2,\dots,m\}$ be a bijection. It holds that
$0<z_{(\rho(1),1)}(t_{\rho})<\dots<z_{(\rho(m-1),1)}(t_{\rho}).$
This implies that the number of best responses is lower bounded by
$(m-1)!\geq\left(\frac{1}{2}\sqrt{n}\right)^{\frac{1}{2}m}=n^{\Omega(m)},$
as needed. ∎
Appendix D An Equivalence Between Coverage Functions and “Correlated Maximum” Functions
In this section we prove a generalization to Proposition 4.2 which may be of independent interest. Specifically we show an equivalence between a set function which maps subsets of correlated random variables to their expected maximum value and coverage functions. In this equivalence the representation size of each is polynomial in the representation size of the other, which implies hardness results with respect to one representation imply hardness results with respect to the other. This equivalence is formalized in the following proposition.
Proposition D.1.
Let $\{X_{i}\}_{i\in A}$ be (correlated) random variables which attain values within some set $M$ and have support $\mathcal{R}\subseteq M^{A}$ and probability density function $p:\mathcal{R}\rightarrow[0,1]$. There exists a coverage function $f=(U,\{w_{u}\}_{u\in U},A,h)$ with $|U|\leq|\mathcal{R}|^{2}$ such that $f(S)=\mathbb{E}[\max_{i\in S}X_{i}]$. Moreover, for every coverage function $f=(U,\{w_{u}\}_{u\in U},A,h)$, there exist (correlated) random variables $\{X_{i}\}_{i\in A}$, with support $\mathcal{R}\subseteq\{0,L\}^{A}$, for some $L\geq 0$, such that $f(S)=\mathbb{E}[\max_{i\in S}X_{i}]$ and $|\mathcal{R}|\leq|U|$.
Proof.
Let $\{X_{i}\}_{i\in A}$ be (correlated) random variables which attain values within some set $M$ and have support $\mathcal{R}\subseteq M^{A}$ and probability density function $p:\mathcal{R}\rightarrow[0,1]$. For any $m\in M$ we denote
$\text{prev}(m)=\max\{x\mid x\in M:x<m\},$
which we interpret as 00 when the set is empty.
We now define the coverage function $f=(U,\{w_{u}\}_{u\in U},A,h)$, where $U=\mathcal{R}\times M$, for each $u=(v,m)\in U$ we define $w_{u}=p(v)(m-\text{prev}(m))$, and $h(i)=\{(v,m)\in U\mid v_{i}\geq m\}$. Note that $|M|\leq|\mathcal{R}|$, which means $|U|=|\mathcal{R}|\cdot|M|\leq|\mathcal{R}|^{2}$ as needed. Now,
$\begin{split}\mathbb{E}[\max_{i\in S}X_{i}]&=\sum_{v\in\mathcal{R}}p(v)\max_{i\in S}v_{i}=\sum_{v\in\mathcal{R}}p(v)\sum_{m\leq\max_{i\in S}v_{i}}(m-\text{prev}(m))\\ &=\sum_{v\in\mathcal{R}}\sum_{m\leq\max_{i\in S}v_{i}}p(v)\ (m-\text{prev}(m))=\sum_{v\in\mathcal{R}}\sum_{m\leq\max_{i\in S}v_{i}}w_{(v,m)}\\ &=\sum_{(v,m)\in U}w_{v,m}\mathbb{1}[\exists i\in S.\leavevmode\nobreak\ v_{i}\geq m]=\sum_{(v,m)\in U}w_{v,m}\mathbb{1}[\exists i\in S.\leavevmode\nobreak\ (v,m)\in h(i)]=f(S),\end{split}$
as needed.
In the other direction, let $f=(U,\{w_{u}\}_{u\in U},A,h)$. Denote $L=\sum_{u\in U}w_{u}$. We define our variables $\{X_{i}\}_{i\in A}$ by sampling an element $x\in U$, where each element $u\in U$ is sampled with probability
$Pr[x=u]=w_{u}\frac{1}{L}.$
For any $i\in A$ and $u\in U$ we denote $X_{i}(u)=L\cdot\mathbb{1}[u\in h(i)]$. Our random variables $\{X_{i}\}_{i\in A}$ are defined as $X_{i}=X_{i}(x)$, for any $i\in A$. Now:
$f(S)=\sum_{u\in U}w_{u}\mathbb{1}[\exists i\in S.\leavevmode\nobreak\ u\in h(i)]=\sum_{u\in U}Pr_{x}[x=u]\cdot L\cdot\mathbb{1}[\exists i\in S.X_{i}(u)=L]=\mathbb{E}_{x}[\max_{i\in S}X_{i}].$
Since each $X_{i}$ is a function of $x$, and there are $|U|$ possible values of $x$ we get that $\mathcal{R}$, which is defined as support of $\{X_{i}\}_{i\in A}$, has size $|\mathcal{R}|\leq|U|$, as needed. ∎